OpenStudy (anonymous):

Factor the expression completely. (If the expression is not factorable using integers, enter PRIME.) x4 − x2 − 42

OpenStudy (anonymous):

(x^2 -7)(x^2 + 6)

OpenStudy (id21):

Let's tell that \[x^2=t\] then \[x^4-x^2-42=t^2-t-42\] let's find the roots of \[t^2-t-42=0\] \[t_{1,2}=\frac{1\pm \sqrt{1+168}}{2}=\frac{1\pm 13}{2}=7;-6\] so you can write\[t^2-t-42\] as a factor:\[(t-7)(t+6)\] Remind that \[t=x^2\] put it in factor and you get: \[(x^2-7)(x^2+6)\]